Please help me solve the problem below by addition, substituting and graphing...
4x - 6y = 12
2x + 2y = 6
Thank you for your help

Answer by rchill(405)

Let's use substitution. First, divide both sides of both equations by 2 and we get 2x-3y=6 and x+y=3. Solve for x in the 2nd equation, we get x=3-y. Substitute 3-y for x in the first equation and get: 2(3-y)-3y=6. That expands to 6-2y-3y=6, and that simplifies to 6-5y=6. Solving for y, we get y=0. To find the value of x, substitute 0 for y in x+y=3; that yields x+0=3, or x=3. To check our answer of (0,3), substitute that into the first equation: 2*3-3*0=6, which simplifies to 6=6, a true statement. That means our answer is correct!